A company purchases an asset with a 5-year depreciable life for $75,00 with no expected salvage value.

The company uses straight line depreciation for financial statements and uses double-declining for tax accounting. Assume a tax rate of 34%.

What is the value of the company's deferred tax account at the end of the third year?

Enter your answer as a whole number with no commas and no dollar sign.
Answer #1

Computation of Depreciation under Straight Line Method

Depreciation = (Cost - Salvage Value) / Number of years of life

Depreciation = (7500 - 0) / 5

Depreciation = $1,500 per year

Computation of Depreciation under Double Declining Method

Rate of Depreciation = 2 * (100 / Number of years of life)

Rate of Depreciation = 2 * (100 / 5)

Rate of Depreciation = 40%

Depreciation Chart of Asset till third year

Year Asset Value at the beginning Depreciation
1 $7,500 $3,000
2 $4,500 $1,800
3 $2,700 $1,080

Computation of Deferred Tax Liability

Year SLM Depreciation (A) DDB Depreciation (B) Difference (B - A) Tax on Difference @ 34%
1 1500 3000 1500 510
2 1500 1800 300 102
3 1500 1080 (420) (142.80)
Total 469.20

At the end of the year 3, the balance in the Deferred Tax account is 469